Heat Transfer - Conduction

A wall is made from an inhomogeneous (nonuniform) material for which the thermal conductivity varies through the thickness according k=ax+b, where a and b areconstants. The heat flux is known to be constant. Determine expressions for the temperature gradient and the temperature distribution when the surface at x=0 is attemperature T1.
Answer #1
k = ax + b
dk/dx = a
heat flux is constant
d(kdT/dx) = 0
(dk/dx)(dT/dx) + kd^2T/dx^2 = 0

aT' + kT" = 0

there should be two boundary conditions to solve this differential equation

1 ) at x=0 T =T1
2)??

one more condtion has to be given
other wise we will get an expression with some unknown constant

aT' + kT" = 0
-a/(ax+b) dx = T"/T' dx
integrating on both sides
gradient dT/dx = C1/(ax+b)

T = C1/a ln(ax + b) + T1 - C1/a (ln b)
answered by: duckie